OpenStudy (anonymous):

Write as a logarithm? 4(1/2)^x=8

OpenStudy (anonymous):

\[4(\frac{ 1 }{ 2 })^{x}=8\]

OpenStudy (ybarrap):

$$ \log_2{4(\frac{ 1 }{ 2 })^{x}}=\log _24+x\log_2{\cfrac{1}{2}}=2-x\log_2 2=1-x=\log_2 8=3\\ $$ So, $$ 1-x=3\\ x=-2 $$

OpenStudy (ybarrap):

Make sense?

OpenStudy (ybarrap):

Correction: $$ \log_2{4(\frac{ 1 }{ 2 })^{x}}=\log _24+x\log_2{\cfrac{1}{2}}=2-x\log_2 2=2-x=\log_2 8=3\\ 2-x=3\\ x=2-3=-1 $$

OpenStudy (jdoe0001):

\(\large{ \bf log_{\color{red}{ a}}{\color{blue}{ b}}=y\implies {\color{red}{ a}}^y={\color{blue}{ b}}\\ \quad \\ \quad \\ 4\left(\frac{ 1 }{ 2 }\right)^{x}=8\implies \left(\frac{ 1 }{ 2 }\right)^{x}=\cfrac{8}{4}\implies {\color{red}{ \left(\frac{ 1 }{ 2 }\right)}}^{x}={\color{blue}{ 2}}}\)

OpenStudy (anonymous):

Yes, thank you!